OpenStudy (anonymous):

Please Help!! Write the sum using summation notation, assuming the suggested pattern continues. -9 - 3 + 3 + 9 + ... + 81

OpenStudy (solomonzelman):

Do you need help at first with figuring out what number term is 81 in the sequence ?

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

d= ?

OpenStudy (anonymous):

i know that the first term is -9 and the difference is 6

OpenStudy (solomonzelman):

Okay good for the start

OpenStudy (solomonzelman):

\(\large\color{darkviolet}{ a_n=a_1+d(n-1) }\) \(\large\color{darkviolet}{ a_n=-9+6(80-1) }\) good with this one ?

OpenStudy (anonymous):

a_n=-9+6(81-1)

OpenStudy (solomonzelman):

My bad it is like this, \(\large\color{darkviolet}{ 81=-9+6(n-1) }\)

OpenStudy (solomonzelman):

\(\large\color{darkviolet}{ 81=-9+6(n-1) }\) \(\large\color{darkviolet}{ 90=6(n-1) }\) \(\large\color{darkviolet}{ 15=(n-1) }\) \(\large\color{darkviolet}{ n=14 }\)

OpenStudy (solomonzelman):

What have I found just now ?

OpenStudy (anonymous):

that 81 is the 14th term

OpenStudy (solomonzelman):

Yes, so there are 14 terms and 81 is the 14th.

OpenStudy (solomonzelman):

Can you take an attempt on the sigma ?

OpenStudy (anonymous):

one sec

OpenStudy (solomonzelman):

sure:)

OpenStudy (anonymous):

\[\sum_{n=0}^{15}\left( -9+6n \right)\]

OpenStudy (solomonzelman):

yes, but then put 13 on the top, so that you have 14 numbers 0,1,2,3 .... 12,13. not 15 terms, and certainly not 16 as you wrote.

OpenStudy (anonymous):

i wrote 15 because it starts off with 14 terms but my only answer could either be 15 on top or infinity

OpenStudy (solomonzelman):

\(\large\color{darkviolet}{ 15=(n-1) }\) here is the error \(\large\color{darkviolet}{ 16=n }\) b/c you add 1 to both sides.

OpenStudy (solomonzelman):

\[\sum_{n=0}^{15}~(-9+6n)~~~\rm ~~~~ is~~~correct:)\]

OpenStudy (anonymous):

yay :D, can you help me with a few more?

OpenStudy (solomonzelman):

Yes, I can try :P

OpenStudy (anonymous):

thanks